OpenStudy (anonymous):

A gas occupies 72.0 mL at 25 oC and 198 KPa. What is the volume of this gas at 45 oC and 129 KPa?

OpenStudy (aaronq):

\[\frac{ P _{1} V _{1} }{ T _{1} }=\frac{ P _{2} V _{2} }{ T _{2} }\]

OpenStudy (anonymous):

what about this: Determine the molar mass of an unknown gas if a 1.08 g sample occupies a volume of 59.3 mL at 33 oC and 746 mm Hg.

OpenStudy (aaronq):

use the ideal gas equation, PV=nRT substitute n=mass/molar mass (abbreviated like this "n=m/M") PV=mRT/M isolate M, M=mRT/PV

OpenStudy (anonymous):

so im using M=mRT/PV?

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

and then solve for V?

OpenStudy (aaronq):

solve for M, molar mass

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (aaronq):

no problem